(x^(2)+3)/(x^(2)+8x+15)+(4)/(x+3)=1
Answer by lwsshak3(11628)

(x^(2)+3)/(x^(2)+8x+15)+(4)/(x+3)=1
..
(x^2+3)/(x^2+8x+15)+(4)/(x+3)=1
=(x^2+3)/((x+3)(x+5))+4/(x+3)=1
LCD=(x+3)(x+5)
(x^2+3)+4(x+5)=(x+3)(x+5)
x^2+3+4x+20=x^2+8x+15
4x+23=8x+15
4x=8
x=4